November 28, 2017. I took the EDS detector (in need of repair) off of the TEM yesterday. This morning I started a bake of the column which will take 20 hours. The scope should be ready by tomorrow morning. I’ll send you a followup email to confirm.
November 15, 2017. The FEI engineer will be here on Monday afternoon and probably Tuesday. The scope won’t be available then. I’ll send you updates.
November 15, 2017. Next Thursday and Friday, Nov 23rd and 24th are university holidays. The labs will be closed to all but unrestricted users on those days. Danielle, Justina and Joe will be in and out. Make appointments with them if you need their help.
November 15, 2017. We have had cooling water problems for a few weeks. Since the last one a week ago a few parts aren’t working right. That isn’t a problem because we aren’t using that yet anyway. The scope is ready to be used in conventional TEM mode.
November 15, 2017. The new FEI 200kV Titan Themis S-TEM is installed an operational. We are training users on the conventional TEM operation mode now. The new S-TEM specialist, Dr. Pinaki Mukherjee, will start work in early January 2018.
November 9, 2017. 9:32 am. Scopes are back online except the STEM. Going there now. November 9, 2017. 8:21 am. The power went off in the M&M and ATDC. I’m getting everything back online now. All should be back online by 9 am. I’ll confirm with a 2nd email.